OpenStudy (anonymous):

a ceiling fan is rotating at 0.96 rev/sec. when turned off, it slows uniformly to a stop in 2.4 min. How many revolutions does the fan make in this time? Can anyone help me with this question?

OpenStudy (anonymous):

rev = 0.96 * 2.4 * 60

OpenStudy (anonymous):

the answer is 69..if i calculate ur working steps, i get 138.24.

OpenStudy (anonymous):

@syakamal are u here?

OpenStudy (anonymous):

i think question is incomplete................ there should be -ve acceleration be given

OpenStudy (anonymous):

really? -ve acceleration is when it rotates clockwise right?

OpenStudy (anonymous):

when it turn-off........there is -ve acceleration which make it stop

OpenStudy (anonymous):

okay..but then how the -ve and +ve signs affect the answer?

OpenStudy (anonymous):

in your question you only mention initial speed and time not the acceleration by which it stops

OpenStudy (anonymous):

okay but then how come the anser is 69, not 138? i'm so confused. Please help me

OpenStudy (anonymous):

when there is -ve acceleration time of each revlution will inc by which it takes less revlution..that is the affect of -ve acceleration

OpenStudy (anonymous):

@syakamal are u here?

OpenStudy (anonymous):

i'm here..why?

OpenStudy (anonymous):

read my reply and tell did you understand it or not

OpenStudy (anonymous):

Alhamdulillah!..i got it..The question do says that it slows uniformly and then it stops

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

can u answer one more question, please?

OpenStudy (anonymous):

Alhamdulillah???

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

it means praise to Allah

OpenStudy (anonymous):

If the angular acceleration considered to be constant, we can use \(\theta=\omega_0*t-\frac12 \alpha*t^2\) I ended up with 69.2 rev (I thought you both already got it :) Just trying to help.)

OpenStudy (anonymous):

yeah..thanks imron

OpenStudy (anonymous):

thankx @imron07

OpenStudy (anonymous):

i have to go now but i will come soon

OpenStudy (anonymous):

@uzumakhi @syakamal : did your browser convert my latext code into equation? Mine didn't.... I wonder what's the problem :/

OpenStudy (anonymous):

uzumakhi, here is the question.. Using the result from part a (69.02), find the number of revolutions the fan must make for its speed to decrease from 0.96 rev/s to 0.48 rev/s.

OpenStudy (anonymous):

omega^2=omega0^2-2*alpha*theta omega=angular vel omega0=initial angular vel alpha=angular acceleration (found in first problem)

OpenStudy (anonymous):

i am back............sorry i have to go to have lunch

OpenStudy (anonymous):

its ok..sorry if i'm disturbing u

OpenStudy (anonymous):

you are not disturbing me i love to solve this type of sums

OpenStudy (anonymous):

initially i was saying the question is incomplete......so contact to your teacher and after that we will together solve it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

its better if you close the question and try another one

OpenStudy (anonymous):

thanks for ur help